OpenStudy (anonymous):

Write a proportion for the following problem. Eric is making a scale drawing of the ground floor of his house. If the scale is 2 in = 5 ft, what number of inches will he use for a length of 55 ft? A. 2/5 = 55/x B. 2/55 = x/5 C. 2/5 = x/55

OpenStudy (anonymous):

2/5 =55/x?

OpenStudy (anonymous):

Aha ! we have a Proportion : 2in__________________________5ft x____________________________55ft Got it ?:)

OpenStudy (anonymous):

so 2/5 = x/55?

OpenStudy (anonymous):

Yep !!!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You re welcome !:)